### 23 Mar 2026 v1

# Multivariable Painlevé-II equation: connection formulas for asymptotic solutions

Abstract
It is shown that a generalization of the Painlevé-II equation (P-II) to a system of coupled equations with symmetry breaking terms is integrable. A Lax pair for this system is used to relate the asymptotic behavior of the solutions at different infinities via an asymptotically exact WKB approach. The analysis relies on an exact solution of the quantum mechanical Demkov-Osherov model (DOM). An application to the problem of unstable vacuum decay during a second order phase transition provides precise scaling of the number of excitations, including subdominant contributions.
I Introduction
Theoretical physics relies on a set of special functions, which are usually solutions of differential equations whose properties can be understood in great detail. Thus, throughout the 20th century, a physicist’s standard toolbox included knowledge of solvable linear ordinary second order differential equations. The commonly used property of such equations is the existence of connection formulas that relate the parameters of the asymptotic solutions in different limits.
Towards the end of the 20th century, the connection formulas for asymptotic solutions of certain second order nonlinear differential equations were also derived. For example, the homogeneous P-II
$u^{\prime\prime}(x)=xu(x)-2u(x)^{3},$ (1)
is one of the most frequently encountered nonlinear second order differential equations in theoretical physics. Connection formulas [14, 12] relate the behavior of $u(x)$ in the $x\to\pm\infty$ limits. They were used to write integrals of Painlevé transcendents over $x$ in terms of classical special functions [1, 16, 4]. Applications of such connection formulas have already been numerous, including in ultracold atoms [13, 25], viscous flow [18], spectroscopy [2], quantum phase transitions [29], quantum information [15], plasma physics [34, 21], liquid crystals [9], and stochastic processes [5].
In a seemingly different direction, there is a progress on solvable higher-order linear systems, including the multistate Landau-Zener models [26, 27, 28, 8] and solvable quantum quench models [20, 19, 33, 3, 22, 23, 24], which describe interacting quantum systems of potentially combinatorial complexity under explicitly time-dependent conditions. This raises the question of whether there exist practically interesting higher-order systems of nonlinear differential equations that are integrable and possess simple connection formulas for the asymptotic behavior of their solutions as well.
The present Letter provides exact analytical connection formulas for a multivariable generalization of P-II. This nonlinear system is shown to be related to the Demkov-Osherov model (DOM) [10, 29], which is one of the earliest known solvable generalizations of the quantum mechanical Landau-Zener model to an arbitrary number of interacting states with linearly time-dependent parameters. As the class of solvable multistate Landau-Zener systems is known to be large [7, 6], its relevance to solvability of nonlinear systems suggests that the class of integrable multi-variable generalizations of the Painlevé equations with tractable asymptotic behavior may be as large. Thus, the synergy between the two research directions in mathematical physics may enable analytical treatment of more complex nonlinear problems.
II Definition of the model
Let a system of $n$ coupled nonlinear differential equations depend on $n$ parameters $e_{1}<e_{2}<\ldots<e_{n}$:
$u_{1}^{\prime\prime}(x)$ $=$ $xu_{1}(x)-2u_{1}(x)\sum_{k=1}^{n}u_{n}^{2}(x)+e_{1}u_{1}(x),$
$u_{2}^{\prime\prime}(x)$ $=$ $xu_{2}(x)-2u_{2}(x)\sum_{k=1}^{n}u_{n}^{2}(x)+e_{2}u_{2}(x),$ (2)
$\cdots$
$u_{n}^{\prime\prime}(x)$ $=$ $xu_{n}(x)-2u_{n}(x)\sum_{k=1}^{n}u_{n}^{2}(x)+e_{n}u_{n}(x).$
A shift, $x\rightarrow x-e_{1}$, sets $e_{1}=0$, which will be assumed here.
The system in Eq. (2) is integrable. The asymptotic behavior of the solution as $x\rightarrow\pm\infty$ can be connected by analytic formulas. The first observation that enables this result is that the system in Eq. (2) is obtained as a consistency condition
$\left(\frac{\partial H}{\partial x}-\frac{\partial H_{1}}{\partial t}\right)-i[H,H_{1}]=0,$ (3)
for two $(n+1)$-dimensional Hermitian matrices, $H(t,x)$ and $H_{1}(t,x)$, given explicitly by
$H$ $=$ $A-4tB+2C,\quad H_{1}\!=\!\left[\begin{array}[]{ccccc}t&-u_{1}&\cdots&-u_{n-1}&-u_{n}\\ -u_{1}&-t&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ -u_{n-1}&0&\cdots&-t&0\\ -u_{n}&0&\cdots&0&-t\end{array}\right],$ (9)
where
$A=$
$\left[\begin{array}[]{ccccc}4t^{2}+x-2\sum\limits_{k=1}^{n}u_{k}^{2}(x)&0&\cdots&\cdots&0\\ 0&-(4t^{2}+x+2e_{1})+2u_{1}^{2}&2u_{1}u_{2}&\cdots&2u_{1}u_{n}\\ \vdots&2u_{2}u_{1}&-(4t^{2}+x+2e_{2})+2u_{2}^{2}&\ddots&2u_{2}u_{n}\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&2u_{n}u_{1}(x)&2u_{n}u_{2}&\cdots&-(4t^{2}+x+2e_{n})+2u_{n}^{2}\end{array}\right],$ (15)
$B=\left[\begin{array}[]{cccc}0&u_{1}&\cdots&u_{n}\\ u_{1}&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ u_{n}&0&\cdots&0\end{array}\right],\quad C=\left[\begin{array}[]{cccc}0&-iu_{1}^{\prime}&\cdots&-iu_{3}^{\prime}(x)\\ iu_{1}^{\prime}&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ iu_{n}^{\prime}&0&\cdots&0\end{array}\right],\quad{\rm where}\quad u_{k}\equiv u_{k}(x),\quad u_{k}^{\prime}\equiv\frac{du_{k}(x)}{dx}.$ (24)
Thus, if the functions $u_{k}(x)$, $k=1,\ldots n$, satisfy Eq. (2), then the Hamiltonians $H$ and $H_{1}$ from Eq. (9) satisfy the consistency condition (3). Similar properties have been previously used in studies of the Painlevé [12] and many other nonlinear differential equations of mathematical physics [11]. Therefore, the fact that similar pairs of Hamiltonians can be found for higher-order nonlinear equations is not surprising. However, the consistency conditions do not guarantee that the WKB analysis that they enable leads to analytically tractable equations. The second observation that enables the following results is that such a WKB analysis for the system in Eq. (2) can be performed completely analytically.
III The system of two variables: main results
Derivation of the connection formulas for the entire system in Eq. (2) is somewhat involved. The goal of this brief Letter is, instead, to announce their existence and demonstrate the consequences of the integrability for the simplest nontrivial case of only two variables, $n=2$ in Eq. (2), while leaving many details to future publications. For $n=2$, the system in Eq. (2) reduces to
$u_{1}^{\prime\prime}(x)$ $\!=\!$ $xu_{1}(x)\!-\!2u_{1}(x)\left[u_{1}^{2}(x)\!+\!u_{2}^{2}(x)\right],$ (25)
$u_{2}^{\prime\prime}(x)$ $\!=\!$ $xu_{2}(x)\!-\!2u_{2}(x)\left[u_{1}^{2}(x)\!+\!u_{2}^{2}(x)\right]\!-\!eu_{2}(x),$ (26)
where $e>0$.
III.1 Connection formulas
Standard perturbative analysis [17] fixes the asymptotic behavior of the solution for the system in Eqs. (25) and (26) as $x\rightarrow$-$\infty$ up to four parameters, $\alpha_{1,2}$ and $\varphi_{1,2}$, where $\alpha_{1,2}>0$. Namely,
${\rm for}$ $x\rightarrow-\infty:$
$u_{1}$ $\!=\!$ $\frac{\alpha_{1}}{(-x)^{1/4}}\sin\left[\frac{2}{3}(-x)^{3/2}+\frac{(3\alpha_{1}^{2}+2\alpha_{2}^{2})}{4}\ln(-x)+\varphi_{1}\right],$
$u_{2}$ $\!=\!$ $\frac{\alpha_{2}}{(-x+e)^{1/4}}\sin[\frac{2}{3}(-x+e)^{3/2}+$ (27)
$\quad\quad\quad\quad\quad\frac{(3\alpha_{2}^{2}+2\alpha_{1}^{2})}{4}\ln(-x)+\varphi_{2}\large].$
The constant parameters $\alpha_{1,2}$ and $\varphi_{1,2}$ will be referred to as initial amplitudes and phases, respectively.
As $x\rightarrow+\infty$, the trivial perturbation theory easily fixes only the leading $x$-dependent terms in the oscillation phases, while the sub-leading terms, starting with the terms behaving as $\propto\ln x$, are derived using the WKB approach described in Section IV. The solution is finally, i.e., as $x\rightarrow+\infty$, parametrized by two positive amplitudes, $\rho$ and $A$, two phases, $\phi_{1}$ and $\phi_{2}$, and one sign parameter, $\sigma=\pm 1$:
${\rm For}$ $x\rightarrow+\infty:$
$u_{1}(x)$ $=$ $\sigma\sqrt{\frac{x}{2}}+$
$\sigma\frac{\rho}{(2x)^{1/4}}\cos\left[\frac{2\sqrt{2}}{3}x^{3/2}-\frac{3}{2}\rho^{2}\ln x+\phi_{1}\right],$
$u_{2}(x)$ $=$ $\sigma A\cos\left[\sqrt{e}x-\frac{A^{2}\sqrt{e}}{2}\ln x+\phi_{2}\right].$ (29)
Let me define the following combinations of the final and initial parameters:
$I_{1}\equiv\frac{\rho^{2}}{2},\quad I_{2}\equiv\frac{A^{2}\sqrt{e}}{2},$ (30)
$p_{1}=e^{-\pi\alpha_{1}^{2}},\quad p_{2}\equiv e^{-\pi\alpha_{2}^{2}},$ (31)
including two phases:
$\Phi_{1}$ $\equiv$ $\frac{\pi}{4}+{\rm arg}\Gamma\left(i\frac{\alpha_{1}^{2}}{2}\right)+\varphi_{1}-\frac{3\alpha_{1}^{2}}{2}\ln 2+\frac{\alpha_{2}^{2}}{2}\ln(e/4),$
$\Phi_{2}$ $\equiv$ $\frac{\pi}{4}+{\rm arg}\Gamma\left(i\frac{\alpha_{2}^{2}}{2}\right)+\varphi_{2}-\frac{3\alpha_{2}^{2}}{2}\ln 2+\frac{\alpha_{1}^{2}}{2}\ln(e/4).$
The main result of this article is the following connection formulas that relate the final parameters $\{\sigma,I_{1},\,I_{2},\phi_{1},\phi_{2}\}$ to the set of initial parameters $\{p_{1},p_{2},\Phi_{1},\Phi_{2}\}$ and the parameter of the equation, $e$:
$\sigma$ $=$ ${\rm sign}\left[\sin\left(\Phi_{1}\right)\right],$ (32)
$I_{1}$ $=$ $-\frac{1}{4\pi}\ln\left(1-p_{1}^{2}p_{2}^{2}\large\left|1+\frac{1-p_{1}}{p_{1}}e^{2i\Phi_{1}}+\frac{1-p_{2}}{p_{1}p_{2}}e^{2i\Phi_{2}}\large\right|^{2}\right),$ (33)
$\phi_{1}$ $=$ $-\frac{3\pi}{4}-I_{1}\cdot 7\ln 2+{\rm arg}\Gamma\left(2iI_{1}\right)-{\rm arg}\left(1+\frac{1-p_{1}}{p_{1}}e^{2i\Phi_{1}}+\frac{1-p_{2}}{p_{1}p_{2}}e^{2i\Phi_{2}}\right),$ (34)
$I_{2}$ $=$ $-\frac{1}{\pi}\ln\left(2\sqrt{p_{2}p_{1}(1-p_{1})}\large\left|\sin\left(\Phi_{1}\right)\large\right|\right)-2I_{1},$ (35)
$\phi_{2}$ $=$ $\frac{3\pi}{4}-\frac{2}{3}e^{3/2}-I_{2}\ln(4e^{1/2})+{\rm arg}\left[\Gamma(iI_{2})\right]-{\rm arg}\left(e^{i\Phi_{2}}+e^{-i\Phi_{2}}\left(p_{1}+(1-p_{1})e^{2i\Phi_{1}}\right)\right).$ (36)
Figure 1: Numerical test of Eqs. (33)-(36). Solid curves correspond to theoretical predictions for the dependence on the equation parameter $e\in(0.05,5)$ (a) of the action variables $I_{1}$ (brown) and $I_{2}$ (blue); (b) of the angle-related variables $\sin\phi_{1}$ (brown) and $\sin\phi_{2}$ (blue). Discrete points correspond to results of numerical simulations. Both for analytical predictions and numerical calculations the initial conditions were chosen to be $\alpha_{1}=0.9$, $\alpha_{2}=0.8$, $\varphi_{1}=\pi/2$, $\varphi_{2}=\pi/3$. Simulations were performed in the interval $x\in(-5000,5000)$ with a discretization step $dx=0.00001$. The algorithm is described in [31]. Both theory and numerical simulations agree on that $\sigma=-1$ for all points here.
Figure 2: Numerical test of Eqs. (32)-(36). Solid curves correspond to theoretical predictions for the dependence on the angle $\varphi_{1}\in(0,\pi)$ (a,b) of the action variables $I_{1}$ (brown) and $I_{2}$ (blue), respectively; (c,d) of the angle-related variables $\sin\phi_{1}$ (brown) and $\sin\phi_{2}$ (blue), respectively. Discrete points correspond to results of numerical simulations. Both for analytical predictions and numerical calculations the initial conditions were chosen to be $e=1$, $\alpha_{1}=0.8$, $\alpha_{2}=0.6$, $\varphi_{2}=\pi/2$. Simulations were performed in the interval $x\in(-5000,5000)$ with a discretization step $dx=0.00001$. Inset in (b) shows the test of Eq. (32) for the sign $\sigma$ dependence on $\varphi_{1}$ (purple).
III.2 Tests of connection formulas
The first test for the validity of Eqs. (32)-(36) is the observation that at $\alpha_{2}=0$, Eq. (26) has a trivial solution $u_{2}(x)=0$, which corresponds to $I_{2}=0$ and $u_{1}(x)$ satisfying Eq. (1). As expected, Eqs. (33) and (34) then reduce to the previously known connection formulas for Eq. (1) listed in [12].
More generally, the connection formulas were tested numerically for $x$-evolution with Eqs. (25) and (26). Simulations were performed using the approach described in [31], which treats $x$ as a time variable. The evolution was considered during $x\in(-5000,5000)$ with initial parameters taken to be $O(1)$. After this interval, a small piece of the trajectory was recorded and the final parameters were inferred by the best fit of this piece with Eqs. (III.1) and (29).
The results are shown in Figs. 1 and 2. Since the angles $\phi_{1,2}$ are determined only up to an integer multiple of $2\pi$, to obtain smooth curves, $\sin\phi_{1}$ and $\sin\phi_{2}$ are plotted instead of the angles themselves. Apparently, despite the highly nonlinear and sometimes singular dependence of the final set of parameters on the initial ones, the numerical results appear to be in perfect agreement with the analytical predictions. Many other similar tests (not shown) were performed for different values, e.g., of $\alpha_{1,2}$, with the same conclusion.
Finally, let me comment on corrections to the connection formulas. For large but finite $x$, they appear starting with the terms $\propto 1/\sqrt{x}$, which decay relatively slowly. They can be calculated using the same WKB approach, but the complete study of such terms appears to be difficult due to the large number of diverse relevant contributions. Some of them are singular in the limit $e\rightarrow 0$. This is why the simulations were restricted to $e>0.05$ – otherwise, a considerably larger integration interval is needed in order to achieve a comparable agreement with theoretical predictions.
Numerically, for the range of the tested parameters, with $e>0.05$, there were two specific corrections observed that dominated over other terms of the same order in $x$. These two corrections can be easily included into the connection formulas, leading to noticeable accuracy improvements at large but finite $x$. First, the regular part of $u_{1}(x)$ in Eq. (III.1) is renormalized by interactions as
$\sigma\sqrt{\frac{x}{2}}\rightarrow\sigma\sqrt{\frac{x-2u_{2}(x)^{2}}{2}},$ (37)
where $u_{2}(x)$ is taken from Eq. (29). This correction vanishes as $\propto 1/\sqrt{x}$, but it is only a factor $\sim 1/x^{1/4}$ smaller than the leading oscillatory term in Eq. (III.1). Therefore, its effects are well visible during simulations as a slowly decaying modulation of oscillations in $u_{1}(x)$.
Second, there is a correction to $\phi_{1}$ in Eq. (34) that follows from the renormalization of the energy of the state $|0\rangle$ in the WKB approach applied to the evolution along the path $P_{\infty}$ (see Section IV for details). It can be included in the connection formulas by subtracting an additional term from the right hand side of Eq. (34):
$\phi_{1}\rightarrow\phi_{1}-2\pi I_{2}\sqrt{\frac{e}{2x}}.$ (38)
Although well suppressed, small effects of this correction were visible even for the longest simulations. This is attributed to its growth with $e$ and the existence of singularities that make $I_{2}$ anomalously large at specific initial conditions.
IV Derivation of connection formulas
IV.1 Strategy
Consistency conditions (3) allow one to define a state vector, $|\Psi(t,x)\rangle$, as a solution of the two-time Schrödinger equation [11]
$i\frac{d|\Psi\rangle}{dt}$ $=$ $H(t,x)|\Psi\rangle,$ (39)
$i\frac{d|\Psi\rangle}{dx}$ $=$ $H_{1}(t,x)|\Psi\rangle,$ (40)
where $H$ and $H_{1}$ are $\times3$ matrices given by Eq. (9) for $n=3$. The evolution operator $U_{P}$ along an arbitrary path $P$ in the two-time space $(t,x)$ can be written as a path-ordered exponent:
$U_{P}={\cal T}_{P}e^{-i\int_{P}\{H\,dt+H_{1}\,dx\}},$ (41)
where ${\cal T}_{P}$ is the path ordering operator, such that factors corresponding to earlier points along $P$ appear further to the right inside the product of evolution exponents $e^{-i\{H\,dt+H_{1}\,dx\}}$.
Equation (41) can be written as
$U_{P}={\cal T}_{P}e^{\int_{P}{\bm{A}}\cdot d{\bm{\tau}}},$ (42)
where ${\bm{\tau}}\equiv(t,x)$ is the two-time point and ${\bm{A}}\equiv(-iH,-iH_{1})$ is a non-Abelian field. This field is flat (has zero curvature) [28], so the result of the evolution in Eq. (42) depends only on the endpoints, which are $(-t_{0},-x_{0})$ and $(t_{0},-x_{0})$ in Fig. 3, but does not depend on the choice of the path connecting these points.
Figure 3: An integration path P{P} (dashed arrow) with $x_{0}\rightarrow$-$\infty$ and $t\in(-t_{0},t_{0})$, where $t_{0}\gg x_{0}$, is deformed into the path ${P}_{\infty}$, such that the horizontal segment of $P_{\infty}$ lies at $x_{0}\rightarrow+\infty$ and $t\in(-t_{0},t_{0})$ (dotted arrows). This deformation does not change the evolution operator in Eq. (41), since the initial and final points of the path remain unchanged. The vertical legs of $P_{\infty}$ have $t=\pm t_{0}\rightarrow\pm\infty$, which makes the evolution along them adiabatic.
Thus, the evolution over the path $P$ shown in Fig. 3 can be calculated in two ways. First, one can consider $t$ changing in the interval $t\in(-t_{0},t_{0})$ at fixed negative $x=-x_{0}$. The evolution operator then takes the form
$U_{P}={\cal T}_{t}e^{-i\int_{-t_{0}}^{t_{0}}H(t,-x_{0})\,dt}.$ (43)
Alternatively, $U_{P}$ can be obtained as an evolution operator along the path $P_{\infty}$ in Fig. 3. The vertical legs of $P_{\infty}$ correspond to evolution at fixed $t$. For example, evolution along the left vertical leg is described by the operator
$U_{vl}={\cal T}_{x}e^{-i\int_{-x_{0}}^{x_{0}}H_{1}(-t_{0},x)\,dx},$
where the corner points of the paths are taken as $t_{0},x_{0}\rightarrow\infty$, with the limit $t_{0}\rightarrow\infty$ taken first.
Let $|k\rangle$, where $k=0,1,2$, be the diabatic states, in which basis the Hamiltonians $H$ and $H_{1}$ are written in Eq. (9). For $t_{0}\rightarrow\infty$, these states coincide with the eigenstates of $H_{1}(-t_{0},x)$. The evolution along the left and right legs of $P_{\infty}$ is adiabatic, resulting only in phase factors $e^{\pm 2it_{0}x_{0}}$. Thus, only the horizontal segment of $P_{\infty}$ contributes to the inter-state transition amplitudes.
Both the horizontal part of $P_{\infty}$ and the path $P$, describe unitary quantum evolution over a pseudo-time $t\in(-t_{0},t_{0})$ at large positive and large negative $x$, respectively. Since the operator $H(t,x)$ depends on $u_{1,2}(x)$, the relations between the asymptotic values of $u_{1,2}(x)$ as $x\rightarrow\pm\infty$ are obtained by equating $U_{P}$ to $U_{P_{\infty}}$. Since the inter-state transition amplitudes are determined by the horizontal paths, at either large positive or large negative $x$, both evolution operators can be evaluated using the WKB approach, which becomes exact in the limits $x\rightarrow\pm\infty$.
IV.2 WKB analysis in the limit $x\rightarrow$-$\infty$
In this case, there are two resonant regions for $t\in(-t_{0},t_{0})$ near $t=\pm\sqrt{|x|}/2,$ where all nonadiabatic transitions occur. For large $|x|$, these resonances are well separated. By rescaling time
$t=\tau\sqrt{|x|},$ (44)
the Schrödinger equation (39) is transformed to $i\frac{d}{d\tau}|\Psi\rangle=H(\tau)|\Psi\rangle$, where
$\small{H(\tau)=\left[\begin{array}[]{ccc}|x|^{3/2}(4\tau^{2}-1)-2\sqrt{x}(u_{1}^{2}(x)+u_{2}^{2}(x))&-4|x|\tau u_{1}-2i\sqrt{|x|}u_{1}^{\prime}&-4|x|\tau u_{2}-2i\sqrt{|x|}u_{2}^{\prime}\\ -4|x|\tau u_{1}+2i\sqrt{|x|}u_{1}^{\prime}&-|x|^{3/2}(4\tau^{2}-1)+2\sqrt{|x|}u_{1}(x)^{2}&2\sqrt{|x|}u_{1}(x)u_{2}(x)\\ -4|x|\tau u_{2}+2i\sqrt{|x|}u_{2}^{\prime}&2\sqrt{|x|}u_{1}(x)u_{2}(x)&-|x|^{3/2}(4\tau^{2}-1)+2\sqrt{|x|}(e+u_{2}(x)^{2})\end{array}\right].}$ (45)
Figure 4: (a) Eigenvalues of the Hamiltonian (45) as functions of $t=\sqrt{|x|}\tau$ at large negative $x$ (blue curves). Here, x=16x=-16; $\alpha_{1}=\alpha_{2}=0.1$, $\varphi_{1}=\pi/3$, $\varphi_{2}=3\pi/4$. The values of $u_{1}(x)$ and $u_{2}(x)$ are approximated by asymptotic formulas in Eq. (III.1). The Hamiltonian (45) is written in the basis of diabatic states. The labels 0, 1, 20,\,1,\,2 define the convention for the diabatic state indices. The diabatic states coincide asymptotically with the eigenstates as $t\rightarrow\pm\infty$. The red and green arrows show two semiclassical trajectories that originate on diabatic level 0 as $t\rightarrow$-$\infty$ and terminate at level 1 as $t\rightarrow+\infty$. The dashed circle encloses the behavior of the energy levels near the pseudo-time point $t=\sqrt{|x|}/2$. (b) Near $t=\sqrt{|x|}/2$, the dynamics is described by the exactly solvable DOM for three levels that cross linearly at fixed couplings $g_{1}^{+}$ and $g_{2}^{+}$. The diagram of diabatic levels shows only the time dependence of diagonal elements of the Hamiltonian, which in the DOM are straight lines in the time-energy plot. (c) Eigenvalues of the Hamiltonian $H(t,x)+x^{3/2}(4\tau^{2}+1)\hat{1}$, as functions of $t=\sqrt{x}\tau$ at large positive $x$ (blue). Here, $x=10$, $A=0.2$, $\rho=0.12$, $e=1.2$. A term proportional to $\hat{1}$ was added to the Hamiltonian to expose the spectrum better. The values of $u_{1}(x)$ and $u_{2}(x)$ are approximated by asymptotic formulas in Eqs. (III.1) and (29), respectively. The red arrows show the unique path that connects diabatic levels 0 and $1$. At large positive $x$, it passes through two avoided crossing points near times $t_{0}$ and $t_{+}$.
The time-dependent spectrum of this Hamiltonian for large negative $x$ is shown in Fig. 4(a). Near $\tau=\pm 1/2$, additional simplifications follow. Let
$g_{1,2}^{-}\equiv 2(|x|u_{1,2}-i\sqrt{|x|}u_{1,2}^{\prime}).$
After shifting the time variable to $\tau_{m}=\tau+1/2$, the Hamiltonian in the vicinity of $\tau_{m}=0$ reduces to the three-state DOM Hamiltonian with one level crossing linearly two parallel levels:
$H_{-1/2}(\tau_{m})=\left[\begin{array}[]{ccc}-4|x|^{3/2}\tau_{m}&g_{1}^{-}&g_{2}^{-}\\ (g_{1}^{-})^{*}&4|x|^{3/2}\tau_{m}&0\\ (g_{2}^{-})^{*}&0&4|x|^{3/2}\tau_{m}+2\sqrt{|x|}e\end{array}\right].$ (46)
Similarly, consider the time interval near $\tau=1/2$ and keep only the relevant $x$-dependent terms. Introduce the couplings
$g_{1,2}^{+}=-2(|x|u_{1,2}+i\sqrt{|x|}u_{1,2}^{\prime}).$
After a shift of time as $\tau_{p}=\tau-1/2$, the Hamiltonian near $\tau_{p}=0$ is again the DOM:
$H_{1/2}(\tau_{p})=\left[\begin{array}[]{ccc}4|x|^{3/2}\tau_{p}&g_{1}^{+}&g_{2}^{+}\\ (g_{1}^{+})^{*}&-4|x|^{3/2}\tau_{p}&0\\ (g_{2}^{+})^{*}&0&-4|x|^{3/2}\tau_{p}+2\sqrt{|x|}e\end{array}\right].$ (47)
The diabatic level diagram for $H_{1/2}$ is shown in Fig. 4(b). The two parallel levels may be close to one another, so that all three states interact simultaneously near the avoided crossings. Nevertheless, this model is solvable analytically.
The DOM exact solution for the scattering amplitudes coincides with the prediction of the independent crossing approximation, in which each encountered diabatic level crossing is treated as an independent two-state Landau-Zener transition between the crossing energy levels, followed by adiabatic evolution between successive level crossings [29]. Therefore, the entire calculation of the elements of the evolution matrix $U_{P}$ reduces to identifying semiclassical trajectories connecting the initial and final states and then finding the corresponding quantum amplitudes using the independent crossing approximation. An example of two semiclassical trajectories connecting the initial state $|0\rangle$ to the final state $|1\rangle$ is shown by the red and green arrows in Fig. 4(a). The transition amplitudes between diabatic states are obtained using the scattering matrix for the Landau-Zener model (listed, e.g., in Chapter 5 of Ref. [29]) to describe transitions between pairs of states at any diabatic level intersection. Between such pairwise scatterings, the system undergoes adiabatic evolution with energies that can be estimated within second-order quantum perturbation theory in the small parameter $1/|x|$.
This WKB approach is equivalent to the textbook independent crossing approximation, which is widely used in theoretical physics and described, e.g., in Section 5.4 of [29]. This version of WKB was used in the derivation of transition amplitudes in multistate Landau-Zener models [7]. The advantage of this approach over the WKB analysis commonly used in studies of Painlevé equations [12] is that the pseudo-time $t$ is never treated as a complex variable, so there is no need to analyze the full Stokes phenomenon to obtain the evolution matrix $U_{P}$.
In the present case, the full matrix $U_{P}$ is not required. It is sufficient to assume that initially the system is in state $|0\rangle$, which has the highest energy as $t\rightarrow$-$\infty$, thus leading to the derivation of transition amplitudes $U_{P,00}$, $U_{P,10}$, and $U_{P,20}$.
IV.3 WKB analysis as $x\rightarrow+\infty$
This limit is simpler because, as illustrated in Fig. 4(c), the level crossing pattern leads only to three elementary avoided crossings between pairs of levels. These crossings are well separated both in energy and in time $t$. For example, the times of diabatic level crossings are given by $t_{0}=0$ and $t_{\pm}\approx\pm x/(4\sqrt{e})$. Therefore, each crossing is described by the scattering matrix for the two-state Landau-Zener model. Moreover, for the initial conditions with only level 0 populated, the avoided crossing at $t_{-}$ is irrelevant, so only a single trajectory connects the initial state $|0\rangle$ to any final diabatic state. Therefore, path interference does not occur. Figure 4(c) illustrates this property by showing the only semiclassical trajectory connecting levels 0 and $1$.
To calculate the scattering amplitudes using the asymptotically exact WKB, the functions $u_{1,2}(x)$ as $x\rightarrow+\infty$ were assumed to take the form given in Eqs. (III.1) and (29) with unknown parameters, including at logarithmic phase contributions. Near each avoided crossing, the couplings were expressed in terms of these unknown parameters, and the scattering amplitudes were estimated using the known scattering matrix for the Landau-Zener model.
After obtaining the amplitudes of $U_{P,k0}$, where $k=0,1,2$, separately for $P$ and $P_{\infty}$, the WKB results were compared order by order in powers of $x$ and $t_{0}$, down to $O(1)$ contributions to the phase and amplitude. Terms that depended on $t_{0}$ and on high powers of $x$ on both sides then canceled, leaving the connection formulas written in Eqs. (32)-(36) and fixing the logarithmic $x$-dependence in Eqs. (III.1) and (29).
V Excitations after unstable vacuum decay
The system in Eqs. (25) and (26) was previously introduced in [31, 30], where it described a passage through a phase transition in scalar field theory. It can be interpreted as Hamiltonian dynamics, in which $x$ is associated with physical time, $t$ (not to be confused with the pseudo-time introduced in the previous section), and $u_{1}(t)$ and $u_{2}(t)$ are coordinates of a particle with two degrees of freedom. Let ${\bm{X}}=(u_{1},u_{2})$, and the corresponding time-dependent Hamiltonian is
${\cal H}(t)=\frac{{\bm{P}}^{2}}{2}-t\frac{{\bm{X}}^{2}}{2}+\frac{{\bm{X}}^{4}}{2}+\frac{eu_{2}^{2}}{2},$ (48)
where ${\bm{P}}=(P_{1},P_{2})$ is the vector of momenta canonically conjugated to $u_{1}$ and $u_{2}$, and where ${\bm{X}}^{2}\equiv u_{1}^{2}+u_{2}^{2}$. The system in Eqs. (25) and (26) follows from Newton’s equations of motion with the Hamiltonian (48).
In the limits $t\rightarrow\pm\infty$, the potential energy for this classical motion is that of a harmonic oscillator, so the amplitudes $\alpha_{1,2}$ are related to the asymptotically conserved, as $t\rightarrow$-$\infty$, adiabatic invariants [34, 29].
${\cal J}_{1,2}\equiv\frac{\alpha_{1,2}^{2}}{2}.$
Similarly, $I_{1}$ and $I_{2}$ from Eq. (30) are identified as asymptotically conserved adiabatic invariants as $t\rightarrow+\infty$. Then, $\varphi_{1,2}$ and $\phi_{1,2}$ are the angles canonically conjugated to, respectively, ${\cal J}_{1,2}$ and $I_{1,2}$.
According to [31, 30], this classical dynamics appears as a saddle-point solution for a quantum field theory describing a time-dependent passage through a second-order quantum phase transition. The difference between ${\cal J}_{1,2}$ and $I_{1,2}$ captures inevitable nonadiabatic effects as the classical phase space trajectory passes through a separatrix. Physically, this event corresponds to crossing the critical point of a phase transition in the corresponding field theory [30]. Semiclassically, the actions are related to the numbers of produced quasiparticles: $I_{1}/\hbar$ for the number of Higgs bosons and $I_{2}/\hbar$ for the number of Goldstone bosons. The latter acquire a small mass due to the symmetry breaking term $\propto e$ in Eq. (48). The analysis in [31] did not provide explicit connection formulas but made several observations that can now be verified.
Thus, a useful feature of Eqs. (33) and (35) is that $I_{1}$ and $I_{2}$ depend on $e$ only through the angles $\Phi_{1}$ and $\Phi_{2}$, which themselves depend linearly on the initial phases $\varphi_{1}$ and $\varphi_{2}$. Therefore, $\langle I_{1}\rangle$ and $\langle I_{2}\rangle$ are independent of $e$, where the brackets denote averaging over a uniform distribution of $\varphi_{1}$ and $\varphi_{2}$. This invariance of the averaged actions was conjectured in [31] based on less rigorous analytical arguments and numerical simulations, and is now confirmed.
Next, the solution in Eqs. (32)-(36) exhibits asymmetry amplification, which was the main finding in [31]. No matter how small the symmetry breaking parameter $e$ is, the dynamics for $u_{1}(t)$ and $u_{2}(t)$ strongly differ from one another as $t\rightarrow+\infty$. While $u_{1}(t)$ has a monotonically growing component, the asymptotic solution for $u_{2}(t)$ oscillates around zero. Note that if $e$ changes sign, the asymptotic solutions in Eqs. (III.1) and (29) are obtained by switching indices $1$ and $2$. The solution at $e=0$ reduces to the standard P-II solution, which is unstable in the present context.
The study in [31] also addressed the following question: Given that initially the system is nearly at the quantum vacuum state, which corresponds, in dimensionless variables here, to
${\cal J}_{1}={\cal J}_{2}\equiv{\cal J}\ll 1,$ (49)
what are the values of the adiabatic invariants as $t\rightarrow+\infty$? Averaging over the initial phases $\varphi_{1}$ and $\varphi_{2}$ was required.
Using that $\int_{0}^{2\pi}\ln(2|\sin(\Phi)|)\,d\Phi=0$, averaging Eq. (36) over $\varphi_{1,2}$ at conditions in Eq. (49) gives that
$\langle I_{2}\rangle+2\langle I_{1}\rangle\approx\frac{1}{2\pi}\ln\left(\frac{1}{2\pi\cal J}\right).$ (50)
Similarly, Eq. (33) with Eq. (49) lead to
$\langle I_{1}\rangle\approx\frac{1}{4\pi}\left[\ln\left(\frac{1}{2\pi{\cal J}}\right)-c_{1}\right],$ (51)
where
$c_{1}$ $=$ $\frac{1}{\pi^{2}}\iint_{0}^{\pi}\ln[4-2\cos(2\Phi_{1})$- (52)
$2\cos(2\Phi_{2})]\,d\Phi_{1}d\Phi_{2}\approx 1.166.$
Substituting Eqs. (51) and (52) into Eq. (50) leads also to
$\langle I_{2}\rangle\approx\frac{c_{2}}{2\pi},\quad c_{2}=c_{1}\approx 1.166.$ (53)
Expressions for $\langle I_{1}\rangle$ and $\langle I_{2}\rangle$ in Eqs. (51) and (53) were conjectured in [31] for ${\cal J}\ll 1$ without an analytical derivation of $c_{1,2}$, while numerical simulations in [31] estimated that $c_{1}\approx 1.14$ and $c_{2}\approx 1.19$, with an uncertainty in the last significant digit. This agrees well with the analytically derived values of these coefficients in Eq. (53). Curiously, the value of the double phase integral in Eq. (52), $\approx 1.166$, is the famous square-lattice spanning tree constant [32].
VI Conclusion
When a system of nonlinear equations emerges from consistency conditions (3) for matrix operators, the nonlinear problem is effectively described by a linear one. The difficulty for higher-order systems of nonlinear equations often lies in the lack of a fully analytical solution of the associated linear model, even when WKB methods are applicable. Nevertheless, the class of known fully solvable multistate linear systems has been growing recently. Any of the newly solved linear systems may emerge in the WKB analysis enabled by the consistency conditions for a certain nonlinear system. The present article provides an example of such a situation, leading to analytical connection formulas for asymptotic solutions of two coupled P-II equations with a minimally broken symmetry.
Acknowledgements.
Author thanks Prof. Alexander Its for useful discussions. This work was carried out under the auspices of the U.S. DoE through the Los Alamos National Laboratory, operated by Triad National Security, LLC (Contract No. 892333218NCA000001).